OpenStudy (anonymous):

Jeff needs to buy food and medication for six cats and one dog. The cost of food and medication for a cat is $5 less than two-thirds of the cost of food and medication for a dog. If Jeff spends a total of $195, the amount he spent on food and medication for the dog is $?

OpenStudy (anonymous):

anybody?:/

OpenStudy (shinalcantara):

let 'x' be the amount for cat's food and medication 'y' be the amount for dog's food and medication since the cost of food and medication for a cat is $5 less than two-thirds of the cost of food and medication for a dog then you'll have the equation as: x = 2/3y - 5 ------equation 1 knowing that the total is 195 and there are 6 cats and a dog then, 6x + y = 195 -----equation 2 substitute equation 1 to equation 2 6x + y = 195 6(2/3y - 5) + y = 195 distribute 6 to the quantity 4y - 30 + y = 195 4y + y = 195 + 30 5y = 225 y = $45

OpenStudy (anonymous):

tHANKU! SO MUCH @shinalcantara UR BOSS :)

OpenStudy (shinalcantara):

yw :)